# Learning parameter dependence for Fourier-based option pricing with tensor trains

### v6

@@ VIII Acknowledgements
IX Author contributions statement
R.S. conceived the project, and R.S. and H.T. wrote the code with help from K.M. and performed numerical simulations. All authors actively discussed the results and contributed to the preparation of the manuscript.
X Appendix
X.1 How to set the tolerance $\epsilon^{\phi,\hat{v}}_{\mathrm{SVD}}$
Figure 4: The maximum error $e_{\mathrm{TT}}$ when applying SVD with various $\epsilon^{\phi,\hat{v}}_{\mathrm{SVD}}$ values to the TTs under the parameter settings of $d=10,\vec{\sigma},\eta=\text{auto}$.
Figure 4 shows the maximum relative error $e_{\mathrm{TT}}$ when applying SVD with varying $\epsilon_{\mathrm{SVD}}$ to the TT with parameters for $d=10,~{}\vec{\sigma}$, obtained through TCI. Here, we fixed the tolerance of SVD $\epsilon^{\tilde{V}}_{\mathrm{SVD}}$ to $10^{-9}$. The tolerance $\epsilon^{\phi,\hat{v}}_{\mathrm{SVD}}$ for SVD was chosen to maintain $e_{\mathrm{TT}}$ smaller than $e_{\mathrm{MC},10^{5}}$. From FIG. 4, it can be seen that $e_{\mathrm{TT}}$ increases sharply between $\epsilon_{\mathrm{SVD}}=10^{-6}$ and $10^{-5}$, suggesting that setting $\epsilon_{\text{SVD}}$ around $5.0\times 10^{-6}$ is appropriate for keeping $e_{\mathrm{TT}}$ smaller than $e_{\mathrm{MC},10^{5}}$. By compressing the bond dimension with SVD, the computational time for the contraction of these optimized TTs also can be kept low.
From the fact that we can keep the maximum relative error $e_{\mathrm{TT}}$ small through SVD, it is suggested that the TTs obtained by TCI contain redundant information. By using SVD to get an optimal approximation in terms of the Frobenius norm, the reduced redundant information could be effectively removed. In addition, it is surprising that in the analysis with $\epsilon_{\text{SVD}}=10^{-6}$, the maximum relative error decreased compared to before compression by SVD. We consider that the error contained in TTs obtained by TCI is eliminated through SVD by chance. We expect that this phenomenon does not occur generally, and in fact, it did not occur for other asset numbers or parameters.

### v7

@@ Acknowledgements
Appendix A How to set the tolerance $\epsilon^{\phi,\hat{v}}_{\mathrm{SVD}}$
Figure 4: The mean absolute error $e_{\mathrm{TT}}$ when applying SVD with various $\epsilon^{\phi,\hat{v}}_{\mathrm{SVD}}$ values to the TTs when incorporating the $\vec{\sigma}$ dependence at $d=10$.
Figure 4 shows the maximum relative error $e_{\mathrm{TT}}$ when applying SVD with varying $\epsilon^{\phi,\hat{v}}_{\mathrm{SVD}}$ to the TT when incorporating $\vec{\sigma}$ dependence at $d=10$ obtained through TCI. Here, we fixed the tolerance of SVD $\epsilon^{\tilde{V}}_{\mathrm{SVD}}$ to $10^{-6}$. The tolerance $\epsilon^{\phi,\hat{v}}_{\mathrm{SVD}}$ for SVD was chosen to maintain $e_{\mathrm{TT}}$ smaller than $e_{\mathrm{MC}}$. From FIG. 4, it can be seen that $e_{\mathrm{TT}}$ increases sharply between $\epsilon_{\mathrm{SVD}}=10^{-6}$ and $10^{-2}$, suggesting that setting $\epsilon_{\text{SVD}}$ around $1.0\times 10^{-6}$ is appropriate for keeping $e_{\mathrm{TT}}$ smaller than $e_{\mathrm{MC}}$. By compressing the bond dimension with SVD, the computational time for the contraction of these optimized TTs also can be kept low.
From the fact that we can keep the error $e_{\mathrm{TT}}$ small through SVD, it is suggested that the TTs obtained by TCI contain redundant information. By using SVD to get an optimal approximation in terms of the Frobenius norm, the reduced redundant information could be effectively removed. In addition, it is surprising that in the analysis with $\epsilon^{\phi,\hat{v}}_{\mathrm{SVD}}$ = $10^{-6}$, the $e_{\mathrm{TT}}$ decreased compared to before compression by SVD. We consider that the error contained in TTs obtained by TCI is eliminated through SVD by chance. We expect that this phenomenon does not occur generally, and in fact, it did not occur for other asset numbers or parameters.